OpenStudy (anonymous):

Find all polar coordinates of point P where P = (7, pi/3)

OpenStudy (fozia):

these are polar cordinats dear

OpenStudy (fozia):

okey

OpenStudy (anonymous):

Again? like what we just worked on?

OpenStudy (anonymous):

x = 7cos(pi/3) y = 7sin(pi/3)

OpenStudy (anonymous):

x = 7/2 or 3.5, y = (7sqrt3)/2

OpenStudy (fozia):

here cospi/3 is 0.5 multiply with 7 and then sinpi/3 is 0.866 multiply by 7 gives u reslt

OpenStudy (anonymous):

.5 * 7 = 3.5 I have to get the non-decimal form so i have to use (7sqrt3)/2, but how do i convert that into "all polar coordinates of point P"?

OpenStudy (anonymous):

Choices: (7, pi/3 + 2nπ) or (-7, pi/3 + 2nπ) (7, pi/3 + 2nπ) or (-7, pi/3 + (2n + 1)π) (7, pi/3 + (2n + 1)π) or (-7, pi/3 + 2nπ) (7, pi/3 + nπ) or (-7, pi/3 + nπ) @fozia

OpenStudy (fozia):

thats the thing i wanted to tell yesturday that if u want possible polar cordinates then use the formula (r,angle+2npi)